# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v2

@@ Abstract
Reconstruction-based methods have significantly advanced unsupervised image anomaly detection involving only normal training images. However, it has been proven that modern neural networks generally have a strong reconstruction capacity and often reconstruct both normal and abnormal samples well, thereby failing to spot anomaly regions by checking the reconstruction quality. To prevent well-reconstructed anomalies, one simple but effective strategy is to perturb normal samples and then map perturbed versions to normal ones. Yet it treats each spatial position equally, disregarding the fact that the foreground locations are inherently more important for reconstruction. Motivated by this, we present a simple yet effective reconstruction framework named Attention-Guided Perturbation Network (AGPNet), which learns to add perturbations guided with an attention mask during training. Specifically, it consists of two branches, i.e., a reconstruction branch and an auxiliary attention-based perturbation branch. The reconstruction branch learns to reconstruct normal samples, while the auxiliary one aims to produce attention masks to guide the noise perturbation process for normal samples. By doing so, we are expecting to synthesize hard yet more informative anomalies for training, which enable the reconstruction branch to learn important inherent normal patterns both comprehensively and efficiently. Extensive experiments are conducted on several popular benchmarks covering MVTec-AD, VisA, and MVTec-3D, and show that AGPNet obtains leading anomaly detection results under few-shot, one-class, and multi-class setups.
@@
journal: Computer Vision and Image Understanding
@@
left=1.5cm, right=1.5cm, top=2.5cm, bottom=2.5cm
1 Introduction
Automatically detecting and localizing anomalies in images, i.e., visual anomaly detection (VAD), has extensive applications in manufacturing Bergmann et al. (2019); Defard et al. (2021). However, this task typically encounters the challenge of cold-start Zavrtanik et al. (2021) where gathering normal samples is easy, whereas acquiring anomalous samples is costly or infeasible, which prohibits supervised learning. Therefore, significant effort has been devoted to developing unsupervised anomaly detection (UAD) algorithms Bergmann et al. (2019); You et al. (2022); Bergmann et al. (2020); Zhang et al. (2023); Deng and Li (2022); Schlüter et al. (2022) that learn to model normality distribution with solely normal images during the training process. During inference, images that are far away from the normality distribution are regarded to be anomalous.
Among existing UAD works, reconstruction-based methods Zavrtanik et al. (2022); Liang et al. (2023) hold great promise and have been extensively studied. Generally, they assume that reconstruction networks exposed to normal-only training images can reconstruct normal ones well, but find difficulty in reconstructing abnormal ones. Thus, an anomaly map can be computed by analyzing the reconstruction error between the input and its reconstructed counterpart. As per this assumption, some methods attempt to reconstruct input images with popular generative models, eg, AutoEncoder Gong et al. (2019); Nguyen et al. (2019); Bergmann et al. (2019), GANs Sabokrou et al. (2018); Kim et al. (2020); Yan et al. (2021), or Difusion-based models He et al. (2024); Zhang et al. (2023). However, reconstructing raw pixel values usually fails when normal and anomalous regions share similar values You et al. (2022). Another direction is reconstructing features since they provide more contextual information than raw pixels You et al. (2022); Shi et al. (2021); You et al. (2022). With the availability of general-purpose large pretrained networks, there has been a growing interest in learning a reconstruction network to reconstruct discriminative features from models pretrained on large-scale datasets, e.g., ImageNet Deng et al. (2009).
Despite the outstanding performance, it has been commonly observed You et al. (2022); Zavrtanik et al. (2021); You et al. (2022) that reconstruction-based methods often encounter the “identity shortcut” issue, a phenomenon of reconstructing anomalies well, due to their uncontrollable generalizability. To alleviate this issue, many strategies Haselmann et al. (2018); Zhou et al. (2020); Zavrtanik et al. (2021) have been presented. One simple yet effective strategy is to perturb input to suppress the reconstruction of anomalies (Fig. 1(a)). For instance, AESc Collin and De Vleeschouwer (2021) introduces a stain perturbation model to corrupt the image with an ellipse shape of irregular edges. During training, they simply restore the corruption. SCADN Yan et al. (2021) generates fixed multi-scale striped masks to erase a part of regions in the input image and learn to reconstruct the missing information. DRAEM Zavrtanik et al. (2021) obtains promising results by recovering pseudo-anomaly disturbed normal images for representation. By perturbing the input, the reconstruction network is forced to learn invariant patterns for reconstruction. Evidently, an effective perturbation scheme is critical. However, prior works simply apply fixed or random masks to perturb input images indiscriminately, regardless of their specific contextual and semantic characteristics. In real-world industrial settings with diverse categories and varying shapes and sizes of foreground objects, such strategies easily induce the reconstruction model to overfit the irrelevant background, leading to suboptimal performance.
In this paper, we argue that it is conducive to learning a better reconstruction model by considering sample-specific content and structural priors for perturbation. To this end, we introduce a simple yet effective reconstruction framework named AGPNet (see Fig. 2) for unsupervised anomaly detection. The core idea behind our framework is to perturb the normal input according to the importance of each location during training. This approach enables the reconstruction network to pay more attention to those important areas, thereby learning invariant normal patterns more compactly and efficiently. To achieve this goal, we carefully design an auxiliary branch that generates attention masks for the noise-adding process ((Fig. 1(b))). Specifically, the auxiliary branch combines two important cues to derive the final attention masks. 1) One is the attention maps from the pre-trained feature extraction layers, which provide strong prior on foreground pixels and important localization cues for the reconstruction network. For instance, industrial categories like screws or toothbrushes typically have small foreground areas while textures may have complex structures occupying the whole image. Thus, it would be more efficient and beneficial to pay attention to these important locations. 2) The second cue is the attention maps from the decoder layers of the main branch, which indicates the importance of each location for the reconstruction task. By incorporating this cue, the auxiliary branch dynamically emulates the role of the reconstruction network, thereby preventing training stagnation by perturbing important localizations more aggressively as training progresses. Finally, we perturb the normal input at the main branch with simple Gaussian noise weighted by the final attention masks. This process further helps the framework reason about invariant normal patterns at both low and high levels, ultimately boosting the reconstruction networks to produce discrepancies regarding anomaly samples.
Our framework has several desirable merits. First, it identifies anomalies at high accuracy under both one-class and multi-class settings. By perturbating more aggressively on the important regions, we are in fact attempting to synthesize hard anomalous locations for training, which helps to learn more discriminative and compact boundaries to spot the abnormal patterns. Besides, the attention masks are sample-aware, accommodating images from diverse categories that exhibit different anomaly types. Second, it is more efficient to train. Anomalies in industrial images usually exhibit diverse structures and textures scattered across any foreground locations. Some anomalies can be easily identified by reconstructing them based on proximal pixels with similar colors, while anomalies with irregular structures near the contour or local edge pixels pose a challenge for reconstruction. Masking these critical areas is more conducive and efficient to representation learning for reconstruction. Lastly, it is a concise and effective framework. The whole main branch is just constructed with a pretrained backbone followed by a simple ViT-based decoder, and the auxiliary branch will be removed after training, incurring no computation cost for inference.
[…]
Our contributions are summarized as follows:
@@ 1 Introduction
    We propose to calculate the attention masks based on prior and learnable attention from the reconstruction network. These masks serve to guide the reconstruction network in comprehensively learning normal patterns across diverse industrial images.
@@ 2.1 Unsupervised Visual Anomaly Detection
Due to the uncertain and scarce nature of visual anomalies, a dominant direction in this area is to formulate this problem as an unsupervised paradigm, also known as unsupervised anomaly detection. It assumes that all images in the training set are normal while the test set contains both normal and abnormal images. A great deal of work has been proposed, for more comprehensive surveys, see Liu et al. (2024); Zipfel et al. (2023). Generally, they can be broadly divided into three groups: synthesizing-based, embedding-based, and reconstruction-based methods.
@@
Synthesizing-based methods Li et al. (2021); Schlüter et al. (2022); Zavrtanik et al. (2022, 2021) synthesize anomalies for training. For instance, DRAEM Zavrtanik et al. (2021) attempts to blend predefined texture images with normal samples. CutPaste Li et al. (2021) simply cuts an image patch and randomly pastes it to another image location. These methods heavily rely on the synthesized anomaly quality. However, due to the uncertainty and unpredictability of anomalies, it is hard and even impossible to synthesize all types of real anomalies with high fidelity.
@@
Embedding-based methods Defard et al. (2021); Roth et al. (2022); Liu et al. (2023); Bae et al. (2023) typically work by first utilizing pretrained models to embed normal images and then employing statistical models to model the normal distributions. During testing, samples are regarded as being abnormal if they are far away from the normal distributions. For instance, PaDiM Defard et al. (2021) models embedded patch features with multivariate Gaussian distribution. PatchCore Roth et al. (2022) exploits memory banks to store nominal features. However, they generally demand more computational resources to store the normal embeddings for the training set. In addition, the computing process to identify anomalies during inference is also time-consuming.
@@
Reconstruction-based methods Zavrtanik et al. (2021); Schlüter et al. (2022); Zhou et al. (2020) hold the assumption that the model pretrained only on normal data can reconstruct normal samples while finding difficulty in reconstructing anomalous during inference. However, this assumption sometimes is not satisfied because of the strong and uncontrollable generalization of deep neural networks. Therefore, many strategies have been proposed to address this issue. Some methods try to add prior from the images Zhou et al. (2020); Shi et al. (2021); Xia et al. (2020) as guidance for reconstructions. For instance, P-Net Zhou et al. (2020) proposes to feed the structure features into the reconstruction network. Other methods Collin and De Vleeschouwer (2021); Ye et al. (2020); Liu et al. (2023) frame anomaly detection as the image restoration problem by corrupting the selected content in the images and then learning to restore the corruptions. several methods Li et al. (2020); Zavrtanik et al. (2021) take the idea of image inpainting by first masking the image randomly and then learning to recover it.
@@
Despite promising results, the above methods focus on the “one for one” paradigm, which trains one separate model for each category and thus is more computationally expensive. Recently, some researchers have shifted their attention to another challenging “one for all” paradigm, which only has to train a single for all categories. UniAD You et al. (2022) is the pioneering work for this task. It observses that under this setup the reconstruction network more easily falls into “identical shortcut” problem. They further introduce learnable queries to the transformer network. Later, OmniAL Zhao (2023) proposes a panel-guided anomaly synthesis method to prevent the model from performing identical reconstruction. RLR He et al. (2024a) introduces learnable reference representations to force the framework to learn normal feature patterns explicitly for all categories. MambaAD He et al. (2024b) is the first to integrate the recent Mamba Gu and Dao (2023) architecture into reconstruction-based anomaly detection frameworks, which achieves efficient global contextual modeling while preserving local features.
@@
Following previous works, our work is also reconstruction-based but deals with both “one for one” and “one for all” paradigms. Unlike most existing methods of designing sophisticated architectures, we present a perturbation scheme that could effectively force a simple plain ViT-based reconstruction decoder to learn invariant normal features. Compared with other perturbation schemes treating each spatial location equally, we use attention masks calculated from the model itself to guide the perturbation. By doing so, our reconstruction network can have a better understanding of the intrinsic and important local normal patterns, which help suppress the anomaly reconstructions effectively.
@@ 2.2 Few-shot Visual Anomaly Detection
Recent research has begun to focus on a more challenging task setting: few-shot anomaly detection. It deals with scenarios where only a few training samples are available. RegAD Huang et al. (2022) employs a siamese network architecture to train a registration network, calculating the Mahalanobis distance between the features of the test image and those of a support set containing only normal samples for anomaly detection. With the rise of vision-language models, WinCLIP+ Jeong et al. (2023) introduces CLIP Radford et al. (2021) for few-shot anomaly detection. PromptAD Li et al. (2024) further develops on CLIP, leveraging manually designed and learned anomaly prompts to expand the distance between anomalous and normal features, thereby assisting in anomaly detection.
Our framework can also be effectively extended to few-shot anomaly detection. Compared with existing works, our works does not rely additional memory bank or complex data augmentations. With the simple yet effective attention-based perturbation scheme, our framework is also sample-efficient, obtaining comparable performance to the current leading few-shot anomaly detection methods.
@@ 3 Method
Problem Formulation. Following previous works Zavrtanik et al. (2021); You et al. (2022), we formulate visual anomaly detection (VAD) as a reconstruction problem. Given a training set $\mathcal{X}_{train}=\{I_{normal}^{i}\}_{i=1}^{N}$ comprising only normal samples, our main goal is to train a reconstruction model $f(\cdot)$ to accurately identify both pixel and image-level anomalies in a test set $\mathcal{X}_{test}=\{I^{j}\}_{j=1}^{M}$ containing both normal and abnormal samples. Note that the problem of VAD can be further divided into two settings: one-class and multi-class. The one-class setting involves training individual models with images from each semantic category, while the multi-class setting aims to train a unified model with all images from diverse semantic categories. The latter is more computationally efficient due to the $\mathcal{O}(1)$ task complexity regardless of the semantic category number. Yet it is more challenging due to the diverse distributions of different categories, making it difficult to learn compact representations. Lastly, in addition to conventional unsupervised anomaly detection, our framework can also be versatilly extended to few-shot anomaly detection. In this setup, only a few images are provided to the framework for training.
@@ 3.1.1 Model Overview
As shown in Fig. 2, our framework mainly consists of two branches. The main reconstruction branch aims to reconstruct the inputs, while the auxiliary branch aims to generate attention masks to guide the perturbation process for the inputs. With the attention mask as perturbation guidance, the reconstruction network tends to learn invariant normal patterns at those crucial locations. After training, the framework can be used for anomaly detection directly by comparing the input and output of the decoder directly, with the discrepancy indicating the anomaly locations. Below we describe each component in detail.
3.1.2 Reconstruction Branch

### v3

# Not All Regions Are Equal: Attention-Guided Perturbation Network for Industrial Anomaly Detection

@@ Abstract
In unsupervised image anomaly detection, reconstruction methods aim to train models to capture normal patterns comprehensively for normal data reconstruction. Yet, these models sometimes retain unintended reconstruction capacity for anomalous regions during inference, leading to missed detections. To mitigate this issue, existing works perturb normal samples in a sample-agnostic manner, uniformly adding noise across spatial locations before reconstructing the original. Despite promising results, they disregard the fact that foreground locations are inherently more critical for robust reconstruction. Motivated by this, we present a novel reconstruction framework named Attention-Guided Perturbation Network (AGPNet) for industrial anomaly detection. Its core idea is to add perturbations guided by a sample-aware attention mask to improve the learning of invariant normal patterns at important locations. AGPNet consists of two branches, i.e., a reconstruction branch and an auxiliary attention-based perturbation one. The reconstruction branch learns to reconstruct normal samples, while the auxiliary one aims to produce attention masks to guide the noise perturbation process for normal samples. By perturbing more aggressively at those important regions, we encourage the reconstruction branch to learn inherent normal patterns both comprehensively and robustly. Extensive experiments are conducted on several popular benchmarks covering MVTec-AD, VisA, and MVTec-3D, and show that AGPNet consistently obtains leading anomaly detection performance across a variety of setups, including few-shot, one-class, and multi-class ones.
@@
journal: Neural Networks00footnotetext: Authors contribute equally. Corresponding author.
1 Introduction
Automatically detecting and localizing anomalies in industrial images has extensive applications in manufacturing [1, 2]. However, this task typically encounters the challenge of cold-start [3] where gathering normal samples is easy, whereas acquiring anomalous samples is costly or infeasible. In recent years, significant effort has been devoted to developing unsupervised anomaly detection (UAD) algorithms [4, 5, 6, 7, 8] that learn to model normality distribution with solely normal images. In general, these methodologies can be further divided into three types [9, 10]: Embedding-based [11, 2], Synthesizing-based [3, 6, 12], and Reconstruction-based [13, 14, 15]. While embedding- and synthesis-based methods often rely on intricate designs, reconstruction-based approaches are more straightforward with superior scalability [16] and have thus been extensively studied.
Reconstruction methods currently represent the state-of-the-art. They rely on one core assumption: reconstruction networks exposed to normal data alone can reconstruct normal ones well, but find difficulty in reconstructing abnormal ones. Thus, an anomaly map can be computed by analyzing the reconstruction error of the input. As per this assumption, some methods attempt to reconstruct input images with popular generative models, i.e., AutoEncoder [17, 18, 4], GANs [19, 20, 21], or Difusion-based models [22, 23]. However, reconstructing raw pixel values usually fails when normal and anomalous regions share similar values [24]. Another direction is reconstructing features since they provide more contextual information than raw pixels [24, 25, 26]. With the availability of large pretrained networks, there has been a growing interest in learning to reconstruct discriminative features from models pretrained on large-scale datasets, e.g., ImageNet [27].
The assumption, however, sometimes fails, as networks may reconstruct anomalies unpredictably. To address this, prior work employs input perturbation and then maps them to normal ones to suppress anomaly reconstruction (Figure. 1(a)). For instance, AESc [28] introduces a stain perturbation model to corrupt the image. SCADN [21] generates fixed multi-scale striped masks to erase parts of regions in the input and learn to reconstruct the missing information. DRAEM [3] obtains promising results by recovering pseudo-anomaly disturbed normal images. UniAD [26] perturbs the features uniformly with noise. By perturbing the input, the reconstruction network is forced to learn invariant patterns for reconstruction. However, these works simply apply fixed or random masks to perturb input images/features indiscriminately, regardless of their specific contextual and semantic characteristics. In real-world industrial settings where there exists diverse categories and varying shapes and sizes of objects, such information provides important prior and could be utilized as auxiliary guidance to strengthen normal feature learning.
Motivated by above, in this paper, we propose to learn a reconstruction model by considering sample-specific content and structural priors for perturbation. To this end, we introduce a simple yet effective reconstruction framework named AGPNet (Figure. 2) for unsupervised anomaly detection. The core idea behind our framework is to perturb the normal input according to the importance of each location during training. This approach enables the reconstruction network to pay more attention to those important areas, thereby learning invariant normal patterns more compactly and efficiently. To achieve this goal, we carefully design an auxiliary branch that generates attention masks for the noise-adding process (Figure. 1(b)). Specifically, the auxiliary branch combines two important cues to derive the final attention masks. The first is the attention maps from the pre-trained feature extraction layers, which provide strong prior on foreground pixels and important localization cues for the reconstruction network. For instance, industrial categories like screws or toothbrushes typically have small foreground areas while textures may have complex structures occupying the whole image. Thus, it would be more efficient and beneficial to pay attention to these important locations. The second cue is the attention maps from the momentum distillation [29] of decoder layers from the main branch, which indicate the importance of each location for the reconstruction task. With this cue, the auxiliary branch dynamically emulates the role of the reconstruction network, thereby preventing training stagnation by perturbing important localizations more aggressively as training progresses. Finally, we perturb the normal input at the main branch with simple Gaussian noise weighted by the final attention masks. This process further helps the framework reason about invariant normal patterns at both low and high levels.
Our framework has several desirable merits. First, it identifies anomalies at high accuracy under both one-class and multi-class settings. By perturbating more aggressively on the important regions, we are in fact attempting to synthesize hard anomalous locations for training, which helps to learn more discriminative and compact boundaries to spot the abnormal patterns. Besides, the attention masks are sample-aware, accommodating images from diverse categories that exhibit different anomaly types. Second, it is more efficient to train. Anomalies in industrial images usually exhibit diverse structures and textures scattered across any foreground locations. Some anomalies can be easily identified by reconstructing them based on proximal pixels with similar colors, while anomalies with irregular structures near the contour or local edge pixels pose a challenge for reconstruction. Masking these critical areas is more conducive to representation learning for reconstruction. Lastly, it is a concise and effective framework. The main branch is just constructed with a pretrained backbone followed by a simple ViT-based decoder, and the auxiliary branch will be removed after training, incurring no computation cost for inference.
@@ 1 Introduction
Our contributions can be summarized as follows:
@@
    We propose to calculate the attention masks based on prior and learnable attention from the reconstruction network. These masks serve to guide the reconstruction network in learning normal patterns across diverse industrial images.
@@ 2.1 Unsupervised Visual Anomaly Detection
Due to the uncertain and scarce nature of visual anomalies, a dominant direction in this area is to formulate this problem as an unsupervised paradigm, also known as unsupervised anomaly detection. It assumes that all images in the training set are normal while the test set contains both normal and abnormal images. A great deal of work has been proposed, for more comprehensive surveys, see [30, 31]. Generally, they can be broadly divided into three groups: synthesizing-based, embedding-based, and reconstruction-based methods.
@@
Synthesizing-based methods [32, 33, 34, 3] synthesize anomalies for training. For instance, DRAEM [3] attempts to blend predefined texture images with normal samples. CutPaste [32] simply cuts an image patch and randomly pastes it to another image location. These methods heavily rely on the synthesized anomaly quality. However, due to the uncertainty and unpredictability of anomalies, it is impossible to synthesize all types of real anomalies with high fidelity.
@@
Embedding-based methods [2, 35, 36, 37] typically work by first utilizing pretrained models to embed normal images and then employing statistical models to model the normal distributions. During testing, samples are regarded as being abnormal if they are far away from the normal distributions. For instance, PaDiM [2] models embedded patch features with multivariate Gaussian distribution. PatchCore [35] exploits memory banks to store nominal features. However, they generally demand more computational resources to store the normal embeddings for the training set. In addition, the computing process to identify anomalies during inference is also time-consuming.
@@
Reconstruction-based methods [38, 39, 40] hold the assumption that the model pretrained only on normal data can reconstruct normal samples while finding difficulty in reconstructing anomalous during inference. However, this assumption sometimes is not satisfied because of the strong and uncontrollable generalization of deep neural networks. Therefore, many strategies have been proposed to address this issue. Some methods try to add prior from the images [40, 25, 41] as guidance for reconstructions. For instance, P-Net [40] proposes to feed the structure features into the reconstruction network. Other methods [28, 42, 43] frame anomaly detection as the image restoration problem by corrupting the selected content in the images and then learning to restore the corruptions. several methods [44, 38] take the idea of image inpainting by first masking the image randomly and then learning to recover it.
@@
Despite promising results, the above methods focus on the “one for one” paradigm, which trains one separate model for each category and thus is more computationally expensive. Recently, some researchers have shifted their attention to another challenging “one for all” paradigm, which only has to train a single for all categories. UniAD [26] is the pioneering work for this task. It observses that under this setup the reconstruction network more easily falls into “identical shortcut” problem. They further introduce learnable queries to the transformer network. Later, OmniAL [45] proposes a panel-guided anomaly synthesis method to prevent the model from performing identical reconstruction. RLR [46] introduces learnable reference representations to force the framework to learn normal feature patterns explicitly for all categories. MoeAD [47] is the first to incorporate Mixture of Experts (MoE) technology into reconstruction-based anomaly detection frameworks.
@@
Following previous works, our work is also reconstruction-based but deals with both “one for one” and “one for all” paradigms. Unlike most existing methods of designing sophisticated architectures, we present a perturbation scheme that could effectively force a simple ViT-based reconstruction decoder to learn invariant normal features. Compared with other perturbation schemes treating each spatial location equally, we use attention masks calculated from the model itself to guide the perturbation. By doing so, our reconstruction network can have a better understanding of the intrinsic and important local normal patterns, which help suppress the anomaly reconstructions effectively.
@@ 2.2 Few-shot Visual Anomaly Detection
Recent research has begun to focus on a more challenging task setting: few-shot anomaly detection. It deals with scenarios where only a few training samples are available. RegAD [48] employs a siamese network architecture to train a registration network, calculating the Mahalanobis distance between the features of the test image and those of a support set containing only normal samples for anomaly detection. With the rise of vision-language models, WinCLIP+ [49] introduces CLIP [50] for few-shot anomaly detection. PromptAD [51] further develops on CLIP, leveraging manually designed and learned anomaly prompts to expand the distance between anomalous and normal features, thereby assisting in anomaly detection.
Our framework can also be effectively extended to few-shot anomaly detection. Compared with existing works, our works does not rely additional expensive memory bank. With the simple yet effective attention-based perturbation scheme, our framework is also sample-efficient, obtaining comparable performance to the current leading few-shot anomaly detection methods.
@@ 3 Method
3.1 Problem Formulation
Following previous works [3, 26], we adopt the reconstruction paradigm for visual anomaly detection. Given a training set $\mathcal{X}_{train}=\{I_{normal}^{i}\}_{i=1}^{N}$ comprising only normal samples, our main goal is to train a reconstruction model $f(\cdot)$ to accurately identify both pixel and image-level anomalies in a test set $\mathcal{X}_{test}=\{I^{j}\}_{j=1}^{M}$ containing both normal and abnormal samples. Note that the problem of VAD can be further divided into two settings: one-class and multi-class. The one-class setting involves training individual models with images from each semantic category, while the multi-class setting aims to train a unified model with all images from diverse semantic categories. The latter is more computationally efficient due to the $\mathcal{O}(1)$ task complexity regardless of the semantic category number. Yet it is more challenging due to the diverse distributions of different categories, making it difficult to learn compact representations. Lastly, in addition to conventional unsupervised anomaly detection, our framework can also be versatilly extended to few-shot anomaly detection. In this setup, only a few normal images are provided for training.
@@ 3.2.1 Model Overview
As shown in Figure. 2, our framework mainly consists of two branches. The main reconstruction branch aims to reconstruct the inputs, while the auxiliary branch aims to generate attention masks to guide the perturbation process for the inputs. With the attention mask as perturbation guidance, the reconstruction network tends to learn invariant normal patterns at those crucial locations. After training, the framework can be used for anomaly detection directly by comparing the input and output of the decoder directly, with the discrepancy indicating the anomaly locations. Below we describe each component in detail.
3.2.2 The detail of AGPNet